OpenStudy (anonymous):

Find all solutions to x^4+81=0

OpenStudy (jchick):

Welcome to OpenStudy!

OpenStudy (jchick):

x = +/- 3 or x^4 - 81 = 0 (x^2 + 9)(x^2 - 9) = 0 (x^2 + 9)(x - 3)(x + 3) = 0 x^2 = -9 x = +/- 3i, +/- 3

OpenStudy (mathstudent55):

I thought the equation was x^4 + 81 = 0, not x^4 - 81 = 0.